# Beyond the Black Box: Theory and Mechanism of Large Language Models

### v1

@@ 7.2.2 Safety and Trustworthiness
Transparency. Transparency in the context of LLMs refers to the extent to which a model’s internal representations, decision processes, and outputs can be inspected, understood, and communicated to humans, and is often made concrete through methods for model interpretability [linardatos2020explainable, zhao2024explainability, luo2024understanding, dang2024explainable]. Here, we therefore view transparency mainly through the lens of interpretability techniques that aim to reveal how LLMs encode information and produce predictions, and how such insights can support safer and more reliable deployment. Concretely, existing work often groups interpretability methods into three broad categories: global, local, and mechanistic interpretability [linardatos2020explainable, zhao2024explainability, luo2024understanding]. Global interpretability seeks to characterise what a model has learned and how linguistic or semantic information is organised across layers and components. For example, Hewitt and Manning [hewitt2019structural] introduce structural probes to test whether syntactic dependency trees are encoded as linear structures in contextual word representations. Tenney et al. [tenney2019you] use edge-probing tasks to measure how a wide range of linguistic phenomena are distributed across layers in contextual encoders and transformers. Local interpretability focuses on explaining individual predictions by attributing them to specific input tokens, features, or intermediate activations. Jain and Wallace [jain2019attention] show that standard attention weights can be weakly correlated with gradient-based importance scores and thus are not always faithful post-hoc explanations, while Wiegreffe and Pinter [wiegreffe2019attention] argue that, under appropriate definitions and evaluation protocols, attention distributions can still provide useful evidence for explanations. Sundararajan et al. [sundararajan2017axiomatic] propose Integrated Gradients, an axiomatic attribution method that assigns each input feature a contribution score and has become a common tool for token-level importance analysis in neural NLP and LLM outputs. Mechanistic interpretability goes a step further by attempting to reverse-engineer specific circuits and features inside LLMs. Elhage et al. [elhage2021mathematical] develop a mathematical framework for transformer circuits that characterises the algorithms implemented by small attention-only transformers. Olsson et al. [olsson2022context] identify “induction heads” as attention heads whose learned algorithm underlies a large fraction of in-context learning. Cunningham et al. [cunningham2023sparse] use sparse autoencoders to decompose LLM activations into more interpretable latent features, enabling finer-grained localisation and intervention on model behaviour. Qian et al. [qian2025demystifying] study the reasoning trajectories of large reasoning models from an information-theoretic perspective, and observe a distinctive “MI peaks” phenomenon where the mutual information between intermediate representations and the correct answer suddenly spikes at a few critical generation steps. They further show that these peaks typically align with tokens that express reflection or logical transition, such as “Hmm”, “Wait”, or “Therefore,” which they term thinking tokens.
@@
Hallucination. Hallucination refers to instances where an LLM generates outputs that are plausible yet incorrect, conflicting with the model’s world knowledge or context. Recent theoretical research generally yields negative conclusions regarding the complete elimination of hallucinations. Xu et al. [xu2024hallucination] prove that hallucination is mathematically inevitable for any computable LLM, regardless of model architecture or data, due to the inherent limitations of computability and learnability. This inevitability is further corroborated through various theoretical lens, including inductive biases [wu2024no], language identification [kalavasis2025limits], Bayes-optimal estimators [liu2025hallucinations], and calibration [kalai2024calibrated, kalai2025language].
@@
Regarding the causes of hallucinations, Zhang et al. [zhang2025law] propose the knowledge overshadowing framework, explaining that dominant knowledge suppresses less frequent knowledge during generation. From the perspective of model architecture, Peng et al. [peng2024limitations] argue that transformer architectures have inherent limitations in performing function composition , while Sun et al. [sun2025and] demonstrate that decoder-only transformers act as subsequence embedding models where dominant input subsequences trigger incorrect outputs. Additionally, Kalai et al. [kalai2025language] argue that post-training benchmarks exacerbate hallucinations by penalizing uncertainty, effectively incentivizing models to guess rather than abstain.
@@
In terms of mitigation, Kalavasis et al. [kalavasis2025limits] emphasize the role of negative examples, proving that access to negative feedback allows for consistent generation with breadth. Wu et al. [wu2024no] suggest that if facts are restricted to a concept class of finite VC-dimension, non-hallucinating generation is achievable via an improper learner. Zhang et al. [zhang2025law] propose amplifying overshadowed knowledge via contrastive decoding to mitigate bias. Kalai et al. [kalai2025language] advocate for the modifications of mainstream evaluations to reward appropriate expressions of uncertainty.
@@
Robustness, fairness and privacy. Ensuring the safe and ethical deployment of LLMs requires addressing critical issues categorized under “Safety and Trustworthiness”. Among these, robustness, fairness, and privacy are paramount. For detailed treatments of these concepts within the LLM domain, we refer to comprehensive surveys (e.g., safety [shi2024large], trustworthiness [huang2024survey, huang2024trustllm, liu2023trustworthy], fairness [li2023survey, gallegos2024bias, chu2024fairness], and privacy [yao2024survey, yan2024protecting, das2025security]).
@@
A vast body of literature was dedicated to the theoretical analysis of robustness [muravev2021certified, ruan2021adversarial], fairness [kleinberg2016inherent, liu2019implicit], and privacy [li2017differential, kairouz2015composition] in traditional machine learning, primarily because these concepts were often well-defined and could be formalized using precise mathematical objectives and metrics. However, in the current landscape of LLMs, the definitions of robustness, fairness, and privacy can occasionally be more ambiguous, lacking simple closed-form mathematical representations. Furthermore, evaluating these properties often requires using other LLMs as judges or evaluators, which introduces subjectivity and complexity [guo2023evaluating, li2024llms, gu2024survey]. Despite this, there is still some theoretical work to study the related topics within LLMs. For example, wolf2023fundamental introduce a theoretical framework called behavior expectation bounds to formally investigate the fundamental limitations of robustness in LLMs. The core theoretical conclusion, built on this framework, is that for any undesired behavior that an aligned model exhibits with a small, finite probability, there exists an adversarial prompt (whose length increases with the model’s complexity) that can trigger this behavior with a probability that approaches one as the prompt length increases. This implies a fundamental “alignment impossibility”: any alignment process that attenuates an undesired behavior but does not remove it entirely (i.e., reduces its probability to a non-zero value) cannot be considered safe against adversarial prompting attacks like jailbreaks [yi2024jailbreak]. The framework also suggests that popular alignment techniques, such as RLHF, may make the model more susceptible to being prompted into undesired behaviors.
@@
Resistance to Misuse. The unauthorized or malicious use of LLMs poses significant risks [liu2023trustworthy], which severely erode public trust and destabilize information ecosystems. To combat these harms, the development and deployment of AI-generated text detection tools [guo2023close] like watermarking [kirchenbauer2023watermark] are becoming critical for identifying machine-generated content and ensuring accountability. This method allows the output of proprietary LLMs to be algorithmically identified as synthetic with a negligible impact on text quality. he2024theoretically introduce a unified theoretical framework for watermarking LLMs that jointly optimizes both the watermarking scheme and the detector, revealing a fundamental trade-off between watermark detectability (Type-II error) and text distortion. christ2024undetectable introduce a cryptographically, formally defining it as being computationally infeasible to distinguish watermarked outputs from those of the original model, even with adaptive queries. christ2024provably prove that the watermark is unremovable under the assumption of adversary uncertainty about the high-quality text distribution, establishing a steep quality degradation versus watermark removal trade-off. hu2023unbiased introduce the concept of an unbiased watermark for LLMs, which is provably $n$-shot-undetectable, meaning the watermarked output distribution is identical to the original, thereby guaranteeing no degradation in text quality. li2025robust introduce a method for robust watermark detection called Truncated Goodness-of-Fit test, which models human edits as a sparse mixture distribution problem, and prove Tr-GoF achieves adaptive optimality by reaching the optimal detection boundary of in an asymptotic regime of decreasing watermark signal outperforming existing sum-based methods. li2025statistical introduce a general statistical framework for watermark detection in LLMs based on hypothesis testing using a pivotal statistic, enabling the rigorous evaluation of detection efficiency through class-dependent efficiency (the rate of Type II error decay). hu2024inevitable propose the two reweight framework and provide a no-go theorem, which proves that it is impossible to simultaneously maintain the highest watermark strength and the highest sampling efficiency when the vocabulary size is greater than two.
@@ 7.3.1 Linear Representation Hypothesis
Empirical investigations have extensively investigated the emergence of interpretable linear structures within the activation spaces of LLMs. For instance, gurnee2023language show that LLMs learn linear representations for spatial and temporal dimensions, effectively mapping geography and history across multiple scales. Similarly, marks2023geometry identify a generalized “truth direction” within the model’s geometry, showing that a simple linear probe can consistently distinguish truthful statements across diverse topics and datasets. qian2024towards explore how trustworthiness concepts evolve during the pre-training stage. By applying linear probing technique intermediate checkpoints, they reveal that concepts related to trustworthiness become linearly separable early in the pre-training phase.
Moving beyond observation to theoretical grounding, jiang2024origins argue that the interplay between the next-token prediction objective and the implicit bias of gradient descent naturally compels the formation of these linear representations in high-dimensional settings. Providing a rigorous geometric framework, park2023linear use counterfactual interventions to formalize the LRH in both input and output spaces, and then introduce a “causal inner product” that unifies the geometric treatment of linear probing and model steering, thereby giving these directions a clear causal interpretation. Furthermore, marconato2024all address the universality of these features by establishing an “all-or-none” identifiability theorem, which proves that such linear properties either hold in all or in none of the distributionally equivalent models under specific conditions. li2025task theoretically analyze the efficacy of “Task Arithmetic,” proving that, under suitable assumptions, linear operations like addition and negation can successfully edit knowledge in nonlinear Transformers and even generalize to out-of-domain tasks.
@@ 7.3.2 Failures in LLM Usage
One surprising failure of generalization in auto-regressive LLMs is the reversal curse [berglund2023reversal]. If a model is trained on a fact in one direction (e.g., “A is B”), it will fail to automatically generalize to the reverse direction (“B is A”). This means models struggle with basic logical symmetry and exhibit near-zero accuracy when tested on the reversed fact. Recent theoretical analysis [zhu2024towards] suggests the reversal curse is a consequence of the asymmetry in the effective model weights learned during the standard (stochastic) gradient descent training process for auto-regressive models. Specifically, the increase of weights for the sequence $\text{A}\to\text{B}$ does not necessarily cause a corresponding increase of weights for the sequence $\text{B}\to\text{A}$. This discovery provides a new theoretical framework and solution direction for understanding and improving the logical reasoning ability of LLMs (including the CoT).
Another famous failure is position bias. Position bias in LLMs refers to the tendency of the model to assign disproportionate importance or attention to information based on its location within a long input context. This often manifests as higher weight given to content at the beginning and end of the input. The specific and more dramatic manifestation of this is the “Lost-in-the-Middle” phenomenon [liu2023lost], where the model’s performance significantly degrades when the crucial, relevant information is placed in the middle of a long input context, even when the model’s overall context window is large enough to contain it. [wu2025on] proposes a graph theory framework to analyze position bias in multi-layer Transformers. The research reveals two key insights: causal masking inherently leads to a bias in attention towards the earlier positions of the sequence, as tokens in the deep layers continuously aggregate the context information of earlier tokens; Meanwhile, relative positional encodings introduce the distance attenuation effect to compete and balance with the deviation of the causal mask.
@@ 8 Related Work
LLMs have become a milestone in the development of artificial intelligence. Cutting-edge models are reshaping our paradigm for natural language research [openai2022introducingchatgpt, guo2025deepseek, bai2023qwen, touvron2023llama, team2023gemini, caruccio2024claude]. These systems have transitioned from specialized tools into general-purpose artifacts capable of human-like reasoning and complex problem-solving. The rapid iteration of these models, driven by massive-scale compute and data, has established a new paradigm in AI development where empirical results often outpace foundational understanding.
This accelerated development has left the internal operations of LLMs largely opaque, as the sheer scale of trillions of parameters introduces complexities that defy traditional statistical learning intuitions [scalinglaw1, scalinglaw2]. A primary challenge in the current literature is the emergence of unpredictable behaviors at scale, such as ICL [brown2020language], complex hallucinations [xu2024hallucination], and the distinct “aha moments” [guo2025deepseek] observed during training. While specific studies have pioneered insights into mechanistic interpretability, the existing body of research remains largely fragmented, with theoretical inquiries often isolated from the end-to-end developmental pipeline.

### v2

@@ 7.2.2 Safety and Trustworthiness
Transparency. Transparency in the context of LLMs refers to the extent to which a model’s internal representations, decision processes, and outputs can be inspected, understood, and communicated to humans, and is often made concrete through methods for model interpretability (Linardatos et al., 2020; Zhao et al., 2024a; Luo and Specia, 2024; Dang et al., 2024). Here, we therefore view transparency mainly through the lens of interpretability techniques that aim to reveal how LLMs encode information and produce predictions, and how such insights can support safer and more reliable deployment. Concretely, existing work often groups interpretability methods into three broad categories: global, local, and mechanistic interpretability (Linardatos et al., 2020; Zhao et al., 2024a; Luo and Specia, 2024). Global interpretability seeks to characterise what a model has learned and how linguistic or semantic information is organised across layers and components. For example, Hewitt and Manning (Hewitt and Manning, 2019) introduce structural probes to test whether syntactic dependency trees are encoded as linear structures in contextual word representations. Tenney et al. (2019) use edge-probing tasks to measure how a wide range of linguistic phenomena are distributed across layers in contextual encoders and transformers. Local interpretability focuses on explaining individual predictions by attributing them to specific input tokens, features, or intermediate activations. Jain and Wallace (2019) show that standard attention weights can be weakly correlated with gradient-based importance scores and thus are not always faithful post-hoc explanations, while Wiegreffe and Pinter (2019) argue that, under appropriate definitions and evaluation protocols, attention distributions can still provide useful evidence for explanations. Sundararajan et al. (2017) propose Integrated Gradients, an axiomatic attribution method that assigns each input feature a contribution score and has become a common tool for token-level importance analysis in neural NLP and LLM outputs. Mechanistic interpretability goes a step further by attempting to reverse-engineer specific circuits and features inside LLMs. Elhage et al. (2021) develop a mathematical framework for transformer circuits that characterises the algorithms implemented by small attention-only transformers. Olsson et al. (2022) identify “induction heads” as attention heads whose learned algorithm underlies a large fraction of in-context learning. Cunningham et al. (2023) use sparse autoencoders to decompose LLM activations into more interpretable latent features, enabling finer-grained localisation and intervention on model behavior. Qian et al. (2025) study the reasoning trajectories of large reasoning models from an information-theoretic perspective, and observe a distinctive “MI peaks” phenomenon where the mutual information between intermediate representations and the correct answer suddenly spikes at a few critical generation steps. They further show that these peaks typically align with tokens that express reflection or logical transition, such as “Hmm”, “Wait”, or “Therefore,” which they term thinking tokens.
@@
Hallucination. Hallucination refers to instances where an LLM generates outputs that are plausible yet incorrect, conflicting with the model’s world knowledge or context. Recent theoretical research generally yields negative conclusions regarding the complete elimination of hallucinations. Xu et al. (2024b) prove that hallucination is mathematically inevitable for any computable LLM, regardless of model architecture or data, due to the inherent limitations of computability and learnability. This inevitability is further corroborated through various theoretical lens, including inductive biases (Wu et al., 2024), language identification (Kalavasis et al., 2025), Bayes-optimal estimators (Liu et al., 2025a), and calibration (Kalai and Vempala, 2024; Kalai et al., 2025).
@@
Regarding the causes of hallucinations, Zhang et al. (2025e) propose the knowledge overshadowing framework, explaining that dominant knowledge suppresses less frequent knowledge during generation. From the perspective of model architecture, Peng et al. (2024) argue that transformer architectures have inherent limitations in performing function composition, while Sun et al. (2025) demonstrate that decoder-only transformers act as subsequence embedding models where dominant input subsequences trigger incorrect outputs. Additionally, Kalai et al. (2025) argue that post-training benchmarks exacerbate hallucinations by penalizing uncertainty, effectively incentivizing models to guess rather than abstain.
@@
In terms of mitigation, Kalavasis et al. (2025) emphasize the role of negative examples, proving that access to negative feedback allows for consistent generation with breadth. Wu et al. (2024) suggest that if facts are restricted to a concept class of finite VC-dimension, non-hallucinating generation is achievable via an improper learner. Zhang et al. (2025e) propose amplifying overshadowed knowledge via contrastive decoding to mitigate bias. Kalai et al. (2025) advocate for the modifications of mainstream evaluations to reward appropriate expressions of uncertainty. More recently, Liu et al. (2025b) characterize three types of uncertainty: document scarcity, limited capability, and query ambiguity, and reveal the fact that current LLMs struggle to accurately identify the root cause and solve it, emphasizing the crucial role of uncertainty in LLM hallucination.
@@
Robustness, fairness and privacy. Ensuring the safe and ethical deployment of LLMs requires addressing critical issues categorized under “Safety and Trustworthiness”. Among these, robustness, fairness, and privacy are paramount. For detailed treatments of these concepts within the LLM domain, we refer to comprehensive surveys (e.g., safety (Shi et al., 2024), trustworthiness (Huang et al., 2024a; d; Liu et al., 2023c), fairness (Li et al., 2023b; Gallegos et al., 2024; Chu et al., 2024), and privacy (Yao et al., 2024b; Yan et al., 2024; Das et al., 2025)).
@@
A vast body of literature was dedicated to the theoretical analysis of robustness (Muravev and Petiushko, 2021; Ruan et al., 2021), fairness Kleinberg et al. (2016); Liu et al. (2019), and privacy (Li et al., 2017; Kairouz et al., 2015) in traditional machine learning, primarily because these concepts were often well-defined and could be formalized using precise mathematical objectives and metrics. However, in the current landscape of LLMs, the definitions of robustness, fairness, and privacy can occasionally be more ambiguous, lacking simple closed-form mathematical representations. Furthermore, evaluating these properties often requires using other LLMs as judges or evaluators, which introduces subjectivity and complexity (Guo et al., 2023d; Li et al., 2024a; Gu et al., 2024). Despite this, there is still some theoretical work to study the related topics within LLMs. For example, Wolf et al. (2023) introduce a theoretical framework called behavior expectation bounds to formally investigate the fundamental limitations of robustness in LLMs. The core theoretical conclusion, built on this framework, is that for any undesired behavior that an aligned model exhibits with a small, finite probability, there exists an adversarial prompt (whose length increases with the model’s complexity) that can trigger this behavior with a probability that approaches one as the prompt length increases. This implies a fundamental “alignment impossibility”: any alignment process that attenuates an undesired behavior but does not remove it entirely (i.e., reduces its probability to a non-zero value) cannot be considered safe against adversarial prompting attacks like jailbreaks (Yi et al., 2024). The framework also suggests that popular alignment techniques, such as RLHF, may make the model more susceptible to being prompted into undesired behaviors.
@@
Resistance to Misuse. The unauthorized or malicious use of LLMs poses significant risks (Liu et al., 2023c), which severely erode public trust and destabilize information ecosystems. To combat these harms, the development and deployment of AI-generated text detection tools (Guo et al., 2023a) like watermarking (Kirchenbauer et al., 2023) are becoming critical for identifying machine-generated content and ensuring accountability. This method allows the output of proprietary LLMs to be algorithmically identified as synthetic with a negligible impact on text quality. He et al. (2024a) introduce a unified theoretical framework for watermarking LLMs that jointly optimizes both the watermarking scheme and the detector, revealing a fundamental trade-off between watermark detectability (Type-II error) and text distortion. Christ et al. (2024b) introduce a cryptographically, formally defining it as being computationally infeasible to distinguish watermarked outputs from those of the original model, even with adaptive queries. Christ et al. (2024a) prove that the watermark is unremovable under the assumption of adversary uncertainty about the high-quality text distribution, establishing a steep quality degradation versus watermark removal trade-off. Hu et al. (2023b) introduce the concept of an unbiased watermark for LLMs, which is provably $n$-shot-undetectable, meaning the watermarked output distribution is identical to the original, thereby guaranteeing no degradation in text quality. Li et al. (2025g) introduce a method for robust watermark detection called Truncated Goodness-of-Fit test, which models human edits as a sparse mixture distribution problem, and prove Tr-GoF achieves adaptive optimality by reaching the optimal detection boundary of in an asymptotic regime of decreasing watermark signal outperforming existing sum-based methods. Li et al. (2025f) introduce a general statistical framework for watermark detection in LLMs based on hypothesis testing using a pivotal statistic, enabling the rigorous evaluation of detection efficiency through class-dependent efficiency (the rate of Type II error decay). Hu and Huang (2024) propose the two reweight framework and provide a no-go theorem, which proves that it is impossible to simultaneously maintain the highest watermark strength and the highest sampling efficiency when the vocabulary size is greater than two.
@@ 7.3.1 Linear Representation Hypothesis
Empirical investigations have extensively investigated the emergence of interpretable linear structures within the activation spaces of LLMs. For instance, Gurnee and Tegmark (2023) show that LLMs learn linear representations for spatial and temporal dimensions, effectively mapping geography and history across multiple scales. Similarly, Marks and Tegmark (2023) identify a generalized “truth direction” within the model’s geometry, showing that a simple linear probe can consistently distinguish truthful statements across diverse topics and datasets. Qian et al. (2024) explore how trustworthiness concepts evolve during the pre-training stage. By applying linear probing technique intermediate checkpoints, they reveal that concepts related to trustworthiness become linearly separable early in the pre-training phase.
Moving beyond observation to theoretical grounding, Jiang et al. (2024b) argue that the interplay between the next-token prediction objective and the implicit bias of gradient descent naturally compels the formation of these linear representations in high-dimensional settings. Providing a rigorous geometric framework, Park et al. (2023) use counterfactual interventions to formalize the LRH in both input and output spaces, and then introduce a “causal inner product” that unifies the geometric treatment of linear probing and model steering, thereby giving these directions a clear causal interpretation. Furthermore, Marconato et al. (2024) address the universality of these features by establishing an “all-or-none” identifiability theorem, which proves that such linear properties either hold in all or in none of the distributionally equivalent models under specific conditions. Li et al. (2025b) theoretically analyze the efficacy of “Task Arithmetic,” proving that, under suitable assumptions, linear operations like addition and negation can successfully edit knowledge in nonlinear Transformers and even generalize to out-of-domain tasks.
@@ 7.3.2 Failures in LLM Usage
One surprising failure of generalization in auto-regressive LLMs is the reversal curse (Berglund et al., 2023). If a model is trained on a fact in one direction (e.g., “A is B”), it will fail to automatically generalize to the reverse direction (“B is A”). This means models struggle with basic logical symmetry and exhibit near-zero accuracy when tested on the reversed fact. Recent theoretical analysis (Zhu et al., 2024b) suggests the reversal curse is a consequence of the asymmetry in the effective model weights learned during the standard (stochastic) gradient descent training process for auto-regressive models. Specifically, the increase of weights for the sequence $\text{A}\to\text{B}$ does not necessarily cause a corresponding increase of weights for the sequence $\text{B}\to\text{A}$. This discovery provides a new theoretical framework and solution direction for understanding and improving the logical reasoning ability of LLMs (including the CoT).
Another famous failure is position bias. Position bias in LLMs refers to the tendency of the model to assign disproportionate importance or attention to information based on its location within a long input context. This often manifests as higher weight given to content at the beginning and end of the input. The specific and more dramatic manifestation of this is the “Lost-in-the-Middle” phenomenon (Liu et al., 2023a), where the model’s performance significantly degrades when the crucial, relevant information is placed in the middle of a long input context, even when the model’s overall context window is large enough to contain it. Wu et al. (2025c) propose a graph theory framework to analyze position bias in multi-layer Transformers. The research reveals two key insights: causal masking inherently leads to a bias in attention towards the earlier positions of the sequence, as tokens in the deep layers continuously aggregate the context information of earlier tokens; Meanwhile, relative positional encodings introduce the distance attenuation effect to compete and balance with the deviation of the causal mask.
@@ 8 Related Work
LLMs have become a milestone in the development of artificial intelligence. Cutting-edge models are reshaping our paradigm for natural language research (OpenAI, 2022; Guo et al., 2025; Bai et al., 2023a; Touvron et al., 2023; Team et al., 2023; Caruccio et al., 2024). These systems have transitioned from specialized tools into general-purpose artifacts capable of human-like reasoning and complex problem-solving. The rapid iteration of these models, driven by massive-scale compute and data, has established a new paradigm in AI development where empirical results often outpace foundational understanding.
This accelerated development has left the internal operations of LLMs largely opaque, as the sheer scale of trillions of parameters introduces complexities that defy traditional statistical learning intuitions (Kaplan et al., 2020b; Hoffmann et al., 2022a). A primary challenge in the current literature is the emergence of unpredictable behaviors at scale, such as ICL (Brown et al., 2020), complex hallucinations (Xu et al., 2024b), and the distinct “aha moments” (Guo et al., 2025) observed during training. While specific studies have pioneered insights into mechanistic interpretability, the existing body of research remains largely fragmented, with theoretical inquiries often isolated from the end-to-end developmental pipeline.